# Gravitational Waves from First-Order Phase Transitions Assisted by Temperature-Enhanced Scatterings

### v1

@@ Abstract
Scatterings whose cross sections increase as the cosmic temperature decreases, known as temperature-enhanced scatterings, can have a significant impact on the thermal effective potential of scalar fields responsible for driving cosmological first-order phase transitions. In this work, we systematically investigate how the inclusion of temperature-dependent corrections to the effective potential affects key phase transition parameters, including the nucleation temperature, latent heat release, and inverse duration parameter. These modifications influence both the strength and duration of the phase transition, which in turn determine the properties of the resulting stochastic gravitational wave (GW) background. Employing semi-analytic computational methods, we evaluate the GW spectra generated under these conditions and compare our predictions with the projected sensitivities of forthcoming detectors such as LISA, DECIGO, and BBO. Our analysis demonstrates that temperature-enhanced scatterings can substantially strengthen the phase transition and produce GW signals that lie within the reach of future observational facilities. This study highlights the importance of temperature-dependent microphysical processes in shaping early Universe cosmological signatures.
@@ 1 Introduction
In parallel, first-order phase transitions (FOPTs) in the early Universe have attracted significant interest as potential sources of a stochastic gravitational wave (GW) background [26, 27, 28, 29, 30, 31, 32, 33, 34, 35]. A strongly first-order phase transition proceeds through bubble nucleation, expansion, and collisions, together with plasma processes such as sound waves and turbulence [36, 37, 38, 39, 40, 41, 42, 43, 32, 44]. These mechanisms can generate GW signals within the reach of planned experiments such as LISA [45, 32, 46], DECIGO [47, 48], BBO [49], ET [50], and CE [51].
It is therefore natural to ask whether the same temperature-enhanced scatterings that can drive baryogenesis might also affect the dynamics of a FOPT, potentially strengthening the transition and enhancing its GW signal. Such scatterings modify the thermal effective potential of the scalar field driving the phase transition, thereby altering the critical and nucleation temperatures, the latent heat released, and the duration of the transition. These parameters directly control the amplitude and spectral shape of the resulting GW background [31, 32, 33, 52, 53, 10, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66].
@@
$\Delta V(\phi,T)=c\,T^{p}\phi^{2},$ (1.1)
where $p<0$ captures the growing impact of scatterings as the temperature decreases. We analyze how this term shifts key phase transition parameters, namely the nucleation temperature $T_{n}$, the strength parameter $\alpha$, and the inverse duration parameter $\beta$ [31, 32, 52].
@@
Our results demonstrate that temperature-enhanced scatterings can significantly amplify a FOPT by increasing its strength parameter $\alpha$ and reducing its inverse duration parameter $\beta$, thereby generating an observable GW signal. This provides an intriguing complementarity: the same mechanism that could produce the BAU may also lead to testable predictions for GW experiments. This connection highlights the power of GW cosmology as a probe of physics beyond the Standard Model and motivates further exploration of explicit particle models that realize this scenario [63, 58, 57, 56, 59, 60, 61, 62, 66, 65, 64, 67, 68, 69, 70, 71, 72].
@@ 2.2 Parametrization of the thermal effective potential
where $V_{0}(\phi)$ is the zero-temperature potential and $\Delta V(\phi,T)$ encodes finite-temperature corrections.
@@
The new ingredient is the contribution of temperature-enhanced scatterings, modeled by the additional term 1.1 with $c>0$ and $p<0$. The negative power $p$ captures the fact that scatterings become stronger as the temperature decreases, effectively increasing the thermal mass term for $\phi$ at lower temperatures. This phenomenological form allows us to explore a wide class of possible microscopic realizations in a model-independent way.
@@
At high temperatures, this term is subdominant compared to the usual $a\,T^{2}\phi^{2}$ thermal corrections (with $a>0$), but as the Universe cools, the term $c\,T^{p}\phi^{2}$ can become significant, potentially delaying or strengthening the phase transition.
@@ 2.3 Phase transition parameters
2.4 Example UV Motivation
To illustrate how temperature-enhanced scatterings can arise from an explicit microphysical model, we sketch a minimal scenario involving a light mediator. Suppose the scalar field $\varphi$ responsible for the phase transition couples to a light scalar mediator $S$, which itself interacts with thermal bath fermions $\psi$. The relevant interaction Lagrangian is given by
$\mathcal{L}\supset-\frac{1}{2}m_{S}^{2}S^{2}-y_{S}\,S\,\bar{\psi}\psi-\mu\,S\,\varphi^{2}.$ (2.3)
In this setup, scatterings of the form $\psi\,\varphi\to\psi\,\varphi$ proceed through $t$-channel exchange of the mediator $S$. The corresponding amplitude contains the propagator term $1/(t-m_{S}^{2})$, which grows as the Mandelstam variable $t$ approaches zero at low temperatures. Consequently, the cross section behaves approximately as
$\sigma(\psi\,\varphi\to\psi\,\varphi)\sim\frac{y_{S}^{2}\,\mu^{2}}{(T^{2}+m_{S}^{2})^{2}},$ (2.4)
where the typical momentum transfer is set by the thermal energy, $|t|\sim T^{2}$. For temperatures $T\gg m_{S}$, this yields a scaling $\sigma(T)\propto T^{-4}$, providing a concrete realization of a cross section that grows as the Universe cools.
At finite temperature, such enhanced scatterings contribute an additional correction to the thermal effective potential for $\varphi$ through loop effects. Integrating out the thermal bath degrees of freedom generates an effective term of the form
$\Delta V(\varphi,T)\propto c\,T^{p}\,\varphi^{2},$ (2.5)
with $p<0$, consistent with our phenomenological parametrization in Eq. 1.1.
This minimal example demonstrates that temperature-enhanced scatterings can naturally emerge in models with light mediators and trilinear couplings. Such ingredients are commonplace in many dark sector or extended scalar sector scenarios, supporting the generality of our approach.

### v2

@@ Abstract
Scatterings whose cross sections increase as the cosmic temperature decreases, known as temperature - enhanced scatterings, can have a significant impact on the thermal effective potential of scalar fields responsible for driving cosmological first-order phase transitions. We show that such effects naturally manifest as finite-temperature self-energy corrections to the scalar mass term, leading to an additional contribution of the form $c\,T^{p}\phi^{2}$ in the effective potential. In this work, we systematically investigate how these loop-induced, temperature-dependent corrections affect key phase transition parameters, including the nucleation temperature, latent heat release, and inverse duration parameter. These modifications influence both the strength and duration of the phase transition, which in turn determine the properties of the resulting stochastic gravitational-wave (GW) background. Employing semi-analytic computational methods, we evaluate the GW spectra generated under these conditions and compare our predictions with the projected sensitivities of forthcoming detectors such as LISA, DECIGO, and BBO. Our analysis demonstrates that finite-temperature scattering effects of this kind can substantially strengthen first-order transitions and produce GW signals that lie within the reach of future observational facilities. The results establish a concrete thermal-field-theoretic origin for temperature-dependent modifications of the scalar potential and emphasize their importance in shaping early-Universe cosmological signatures.
Keywords: Gravitational Waves, Cosmological Phase Transitions, Early Universe.
@@ 1 Introduction
In parallel, first-order phase transitions (FOPTs) in the early Universe have attracted significant interest as potential sources of a stochastic gravitational wave (GW) background [26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39]. A strongly first-order phase transition proceeds through bubble nucleation, expansion, and collisions, together with plasma processes such as sound waves and turbulence [40, 41, 42, 43, 44, 45, 46, 47, 33, 48]. These mechanisms can generate GW signals within the reach of planned experiments such as LISA [49, 33, 50], DECIGO [51, 52], BBO [53], ET [54], and CE [55].
It is therefore natural to ask whether the same temperature-enhanced scatterings that can drive baryogenesis might also affect the dynamics of a FOPT, potentially strengthening the transition and enhancing its GW signal. Such scatterings modify the thermal effective potential of the scalar field driving the phase transition, thereby altering the critical and nucleation temperatures, the latent heat released, and the duration of the transition. These parameters directly control the amplitude and spectral shape of the resulting GW background [32, 33, 34, 56, 57, 10, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71].
From the standpoint of finite-temperature quantum field theory, temperature-enhanced scatterings correspond to loop-level corrections to the scalar two-point function. The resulting finite-$T$ self-energy $\Pi_{\phi}(T)$ modifies the thermal mass and hence the curvature of the effective potential near the origin. In the local limit of the one-particle-irreducible effective action, this effect appears as an additional quadratic term in the potential, $\Delta V(\phi,T)\propto\Pi_{\phi}(T)\,\phi^{2}$. Depending on the mediator spectrum and the thermal scaling of the underlying cross sections, the self-energy can behave as $\Pi_{\phi}(T)\!\propto\!T^{p}$ with either positive or negative exponent $p$, thereby motivating the phenomenological parameterization adopted below. This correction is conceptually distinct from the collision terms that enter Boltzmann equations for wall friction or transport—here we focus exclusively on its static contribution to the effective potential.
A temperature-dependent quadratic correction of this form arises naturally in several concrete particle-physics models. In particular, singlet-extended electroweak scenarios (xSM) generate both (i) a tree-level barrier capable of producing a first-order electroweak phase transition, and (ii) loop-induced thermal self-energy corrections that modify the Higgs curvature as $T$ evolves. The combination of Higgs–singlet mixing and thermal dressing of the singlet propagator yields precisely a correction of the structure $\Delta V(\phi,T)=c\,T^{p}\phi^{2}$, with the exponent $p$ determined by the dominant microscopic scattering processes. A representative realization of this mechanism is developed explicitly in Sec. 2.4, where the mapping between the microscopic Lagrangian and the effective parameters $(c,p)$ is derived from the corresponding thermal loop diagrams.
@@
$\Delta V(\phi,T)=c\,T^{p}\phi^{2},$ (1)
where $p<0$ captures the growing impact of scatterings as the temperature decreases. The full finite-temperature potential used in this study thus reads
$V_{\text{eff}}(\phi,T)=V_{0}(\phi)+V_{\text{CW}}(\phi)+V_{T}(\phi,T)+c\,T^{p}\phi^{2},$ (2)
where $V_{0}(\phi)$ denotes the tree-level potential of the model, $V_{\text{CW}}$ the Coleman–Weinberg one-loop term, and $V_{T}$ the usual thermal contribution from particles in equilibrium. Explicit expressions are provided in Sec. 2. We analyze how this term shifts key phase transition parameters, namely the nucleation temperature $T_{n}$, the strength parameter $\alpha$, and the inverse duration parameter $\beta$ [32, 33, 56].
@@
Our results demonstrate that temperature-enhanced scatterings can significantly amplify a FOPT by increasing its strength parameter $\alpha$ and reducing its inverse duration parameter $\beta$, thereby generating an observable GW signal. This provides an intriguing complementarity: the same mechanism that could produce the BAU may also lead to testable predictions for GW experiments. Moreover, by grounding the $T$-dependent correction in explicit finite-temperature self-energy effects, our analysis connects microphysical scattering processes to macroscopic cosmological observables in a self-consistent manner. This connection highlights the power of GW cosmology as a probe of physics beyond the Standard Model and motivates further exploration of explicit particle models that realize this scenario [67, 62, 61, 60, 63, 64, 65, 66, 70, 69, 68, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82].
@@ 2.1 Temperature-enhanced scatterings
It is important to emphasize that a temperature–dependent interaction rate enters the effective potential only through equilibrium self-energy corrections, i.e. the real part of the thermal two-point function. Collision terms appearing in Boltzmann transport equations describe non-equilibrium frictional effects and do not alter the free energy of the system. In our treatment the latter are handled separately through the bubble-wall velocity, while the equilibrium self-energy contribution modifies $V_{\mathrm{eff}}$ via a local $\phi^{2}$ operator.
@@ 2.2 Parametrization of the thermal effective potential
where $V_{0}(\phi)$ is the zero-temperature potential of the model and $\Delta V(\phi,T)$ encodes finite-temperature corrections.
The standard one-loop corrections are included via the Coleman–Weinberg and thermal integrals,
$V_{\rm CW}(\phi)=\sum_{i}\frac{n_{i}}{64\pi^{2}}m_{i}^{4}(\phi)\Big[\ln\frac{m_{i}^{2}(\phi)}{\mu^{2}}-C_{i}\Big],\quad V_{T}(\phi,T)=\sum_{i}\frac{n_{i}T^{4}}{2\pi^{2}}J_{B/F}\!\left(\frac{m_{i}^{2}(\phi)}{T^{2}}\right),$ (5)
where $n_{i}$ counts degrees of freedom, $m_{i}(\phi)$ are field-dependent masses, and $J_{B/F}$ are the standard bosonic/fermionic thermal functions [85].
All thermal functions $J_{B,F}(m^{2}/T^{2})$ are evaluated using the full integral expressions, without invoking any high- or low-temperature series expansions. We follow the numerical prescription of Ref. [86], which directly integrates the exact thermal potentials and therefore remains valid in the domain $m^{2}/T^{2}=\mathcal{O}(1)$ where asymptotic expansions fail. This guarantees that the scattering-induced term $cT^{p}\phi^{2}$ supplements, but does not double count, the equilibrium thermal contributions.
We emphasize that different sectors of the theory are treated in different thermal regimes. The cubic term and tree-level barrier responsible for the first-order phase transition arise from light degrees of freedom (such as gauge bosons and scalars with field-dependent masses $m_{i}(\phi)\lesssim T$ near the transition), for which the standard high-temperature expansion of the thermal functions is valid. In contrast, the mediator field $S$ responsible for the temperature-enhanced scattering contribution is not assumed to be light. Its effect is incorporated solely through its static, thermally dressed propagator entering the scalar self-energy, without invoking a high-temperature expansion for $S$ itself. As a result, the assumptions $m_{i}(\phi)\ll T$ for barrier-generating fields and $m_{S}\gtrsim T$ for the mediator are not in conflict.
2.2.1 Thermal self-energy origin of the $cT^{p}\phi^{2}$ term
From thermal field theory, the leading finite-temperature correction to a scalar field arises from the one-loop self-energy
$\Pi_{\phi}(T)=\mathrm{Re}\,\Sigma_{\phi}(p_{0}=0,\mathbf{p}\!\to\!0;T),$ (6)
which modifies the effective potential as
$\Delta V(\phi,T)\simeq\tfrac{1}{2}\Pi_{\phi}(T)\,\phi^{2}.$ (7)
For heavy mediators one obtains $\Pi_{\phi}(T)\propto+T^{2}$—the usual positive thermal mass. For light $t$-channel mediators, however, the typical momentum transfer scales as $|t|\!\sim\!T^{2}$ and the loop amplitude behaves as $\sim 1/(T^{2}+m_{S}^{2})^{2}$, giving $\Pi_{\phi}(T)\propto T^{p}$ with negative $p$ in the regime $T\gg m_{S}$.
In this local, static limit the contribution is a genuine equilibrium correction to the free energy, and does not reproduce any transport/friction effect that would otherwise enter the Boltzmann equations for bubble-wall propagation. Hence no double counting occurs when the mediator is not already part of the thermal particle content entering $V_{T}(\phi,T)$.
The additional term $cT^{p}\phi^{2}$ represents the equilibrium self-energy contribution of a specific microscopic degree of freedom that is not part of the thermal sum defining $V_{T}(\phi,T)$. Since $V_{T}$ includes only fields with thermal occupation in the plasma, adding $cT^{p}\phi^{2}$ does not duplicate any existing contribution. Non-equilibrium effects such as friction arise only in transport equations and are treated independently in Sec. 3.2, further ensuring that equilibrium and non-equilibrium effects are not mixed.
@@ 2.2.1 Thermal self-energy origin of the $cT^{p}\phi^{2}$ term
The new ingredient is the contribution of temperature-enhanced scatterings, modeled by the additional term 1 with $c>0$ and $p<0$. The negative power $p$ captures the fact that scatterings become stronger as the temperature decreases, effectively increasing the thermal mass term for $\phi$ at lower temperatures. This phenomenological form allows us to explore a wide class of possible microscopic realizations in a model-independent way. From a field-theoretic viewpoint, temperature-enhanced scatterings effectively modify the scalar two-point function through the thermal self-energy $\Pi_{\phi}(T)$. When the relevant interaction cross section scales as $\langle\sigma v\rangle\!\propto\!T^{p}$, the loop integral contributing to $\Pi_{\phi}(T)$ develops an IR enhancement $\Pi_{\phi}(T)\!\sim\!c\,T^{p}$ for $p<0$. This leads to an additional quadratic term $c\,T^{p}\phi^{2}$ in the finite-temperature potential, which parametrizes the dominant low-momentum contribution of such scatterings in a model-independent way.
@@
At high temperatures, this term is subdominant compared to the usual $a\,T^{2}\phi^{2}$ thermal correction (with $a>0$), but as the Universe cools, the term $c\,T^{p}\phi^{2}$ can become significant, potentially delaying or strengthening the phase transition.
Analytic intuition can be gained by expanding near the origin:
$V_{\rm eff}(\phi,T)\simeq\tfrac{1}{2}m_{\rm eff}^{2}(T)\phi^{2}-ET\phi^{3}+\tfrac{1}{4}\lambda\phi^{4},\qquad m_{\rm eff}^{2}(T)=-\mu^{2}+aT^{2}+cT^{p}.$ (8)
For negative $p$, the $cT^{p}$ term grows relative to $aT^{2}$ as $T$ decreases, delaying the sign flip of $m_{\rm eff}^{2}(T)$ and thus increasing supercooling. This lowers $T_{n}$, increases the released latent heat $\Delta\rho$, and therefore enhances the strength parameter $\alpha$ while reducing $\beta/H$, in agreement with our numerical results.
@@
2.2.2 Explicit thermal-QFT example: one-loop derivation of a $cT^{p}\phi^{2}$ term
To demonstrate explicitly how temperature-enhanced scatterings can generate an effective quadratic operator of the form $cT^{p}\phi^{2}$ in the finite-temperature potential, we present a concrete microscopic example and compute the relevant thermal self-energies. The calculation below follows standard finite-temperature field theory (Matsubara) techniques; useful references are [85].
Consider the scalar field $\phi$ (the order parameter) coupled to a mediator scalar $S$ and a fermion $\psi$ in the thermal bath. The interaction Lagrangian is
$\mathcal{L}_{\rm int}\;=\;-\frac{1}{2}m_{S}^{2}S^{2}-\mu\,S\,\phi^{2}-y_{S}\,S\,\bar{\psi}\psi+\mathcal{L}_{\psi,\rm kin},$ (9)
where $\mu$ is a trilinear coupling and $y_{S}$ is a Yukawa coupling between the mediator and the bath fermion.Other UV completions (for example a direct quartic $\lambda_{\phi\chi}\phi^{2}\chi^{2}$ coupling) lead to related but quantitatively different scalings; the present choice cleanly illustrates the IR-enhancement mechanism discussed in the text.
The $\phi$ two-point function receives a contribution from the exchange of the mediator $S$ dressed by fermion loops. In the static, zero-momentum limit the leading contribution to the $\phi$ thermal mass is
$\delta m_{\phi}^{2}(T)\;=\;\mu^{2}\,D_{S}(p_{0}=0,\mathbf{p}\to 0;T),$ (10)
where $D_{S}^{-1}(k_{0},\mathbf{k};T)=k_{0}^{2}+\mathbf{k}^{2}+m_{S}^{2}+\Pi_{S}(k_{0},\mathbf{k};T)$ is the full (thermal) inverse propagator of $S$ and $\Pi_{S}$ is the $S$-self-energy arising from the fermion loop. In the local (infrared) approximation one may evaluate $D_{S}$ at vanishing external momentum; this yields the effective local operator $\tfrac{1}{2}\delta m_{\phi}^{2}(T)\phi^{2}$ in the free-energy.
At one loop the mediator self-energy from the fermion $\psi$ is (Matsubara sum form)
$\Pi_{S}(0,\mathbf{0};T)\;=\;-y_{S}^{2}\,T\sum_{n\in\mathbb{Z}}\int\!\frac{d^{3}\mathbf{k}}{(2\pi)^{3}}\;\mathrm{Tr}\Big[\,G_{\psi}(i\omega_{n},\mathbf{k})\,G_{\psi}(i\omega_{n},\mathbf{k})\,\Big],$ (11)
where $G_{\psi}(i\omega_{n},\mathbf{k})={1}/{(i\omega_{n}\gamma^{0}-\bm{\gamma}\cdot\mathbf{k}-m_{\psi})}$ is the fermion propagator and $\omega_{n}=(2n+1)\pi T$ are fermionic Matsubara frequencies. Performing the Dirac trace and the Matsubara sum leads to the well-known result (for a Dirac fermion) in the high-temperature / massless-fermion limit§§§We note that more general expressions (including finite fermion mass) can be written using the thermal integral representations; the qualitative behaviour relevant for our argument is captured by the displayed limiting form.:
$\Pi_{S}(0,\mathbf{0};T)\;\simeq\;\frac{y_{S}^{2}}{6}\,T^{2}\qquad(m_{\psi}\ll T).$ (12)
This result is textbook material (see e.g. [85]) and corresponds to the thermal screening of the scalar mediator due to the bath fermions.
Using (10) and (12) we obtain
$\delta m_{\phi}^{2}(T)\;=\;\frac{\mu^{2}}{m_{S}^{2}+\Pi_{S}(0,\mathbf{0};T)}\;\simeq\;\frac{\mu^{2}}{m_{S}^{2}+\tfrac{y_{S}^{2}}{6}T^{2}}.$ (13)
Equation (13) is the central, explicit result for this minimal model. It is exact at one-loop if we evaluate $\Pi_{S}$ in the static limit and then perform the algebraic substitution in the dressed propagator. Eq. (13) demonstrates that the $\phi$ thermal mass is controlled by the ratio $\mu^{2}/(m_{S}^{2}+\mathcal{O}(y_{S}^{2}T^{2}))$ and hence exhibits distinct asymptotic behaviours:
  • Mediator-dominated regime: If $m_{S}^{2}\gg y_{S}^{2}T^{2}$, then $\delta m_{\phi}^{2}(T)\simeq\mu^{2}/m_{S}^{2}$ is approximately temperature independent (corresponding to $p\simeq 0$ in our phenomenological parametrization).
  • Thermally-dressed mediator (IR-enhanced) regime: If $y_{S}^{2}T^{2}\gg m_{S}^{2}$, then
    $\delta m_{\phi}^{2}(T)\simeq\frac{6\mu^{2}}{y_{S}^{2}}\,T^{-2},$
    i.e. $\delta m_{\phi}^{2}(T)\propto T^{-2}$ (so p=2p=-2 in the parametrization $cT^{p}$). This negative power arises because the fermion-loop dressing of the mediator produces a thermal screening mass $\Pi_{S}\propto T^{2}$ in the denominator of the dressed propagator, causing the overall $\phi$ mass correction to scale as the inverse of $T^{2}$.
Thus, in this explicit model one obtains a concrete realization of $cT^{p}\phi^{2}$ with $p\leq 0$ depending on the parametric hierarchy between $m_{S}$ and $y_{S}T$. Importantly, the negative power p=2p=-2 occurs naturally in the regime where the mediator is light compared to the thermal scale but acquires a thermal self-energy that grows like $T^{2}$.
Although the derivation above used the dressed-mediator picture, the same physical origin can be viewed from the perspective of $\to2$ scatterings in the thermal bath: $\psi\phi\to\psi\phi$ proceeds dominantly via $t$-channel $S$-exchange with amplitude $\mathcal{M}\sim\mu y_{S}/(t-m_{S}^{2})$. In the thermal average the typical momentum transfer satisfies $|t|\sim T^{2}$, so the squared amplitude scales as $|\mathcal{M}|^{2}\propto\mu^{2}y_{S}^{2}/(T^{2}+m_{S}^{2})^{2}$. Thermally averaging and integrating over phase space yields an effective rate with inverse powers of $T$ in regimes where $m_{S}$ is small; the dressed-mediator self-energy computation above captures this effect in equilibrium as a modification of the static propagator and therefore as a $\phi^{2}$-type contribution to the effective potential. Both viewpoints are consistent: the equilibrium self-energy is the field-theoretic image of forward (coherent) scatterings with the thermal bath.
The example shows that $p$ is model-dependent and can take values $p\leq 0$ depending on the microscopic spectrum and couplings. For the toy model (9) we found $p=0$ (heavy mediator) or p=2p=-2 (light mediator with thermal dressing). Other microscopic scenarios with more infrared-divergent kernels (for example, additional long-range interactions or multiple-step dressing) can produce stronger inverse-temperature scalings in restricted parameter regions; however the dressed-propagator argument above shows that the appearance of a quadratic-in-$\phi$ thermal operator is generic, while the precise exponent $p$ is fixed by the dominant thermal dressing mechanism.
Connection with the effective potential: Working in the Matsubara formalism, the static (zero-frequency, zero-momentum) limit of the one-particle-irreducible two-point function is precisely the second derivative of the free-energy functional with respect to the field and thus corresponds to a quadratic term in $V_{\rm eff}$:
$\Pi_{\phi}(0,0;T)\;=\;\left.\frac{\partial^{2}V_{\rm eff}(\phi,T)}{\partial\phi^{2}}\right|_{\phi=0}.$
The replacement $\tfrac{1}{2}\Pi_{\phi}(T)\phi^{2}$ in $V_{\rm eff}$ is therefore exact at the level of the local, infrared-dominated contribution. This is why the equilibrium self-energy calculation is the appropriate framework to incorporate temperature-dependent scattering effects into the static effective potential used to compute nucleation and tunneling. Non-equilibrium collision terms (Boltzmann integrals) govern transport and friction; they affect wall dynamics but do not replace the equilibrium thermal mass computation above.
Caveats and practical implementation:
  • The relation (13) was derived in the static limit and at one-loop order. It captures the leading equilibrium contribution relevant for the effective potential and nucleation computation. Higher-loop corrections and non-local momentum-dependent pieces can modify the numerical coefficients and produce subleading corrections to the shape of the potential; these effects can be incorporated systematically if required.
  • For numerical work we insert the parametrized form $\delta m_{\phi}^{2}(T)=cT^{p}$ into $V_{\rm eff}$. The constant $c$ and exponent $p$ should be chosen (or matched) according to the microscopic relations such as (13); for the toy model the mapping is $c\simeq 6\mu^{2}/y_{S}^{2}$ and p=2p=-2 in the thermally-dressed regime.
  • The derivation above demonstrates that the extra term is an equilibrium self-energy effect and therefore is not double-counted if the mediator and its thermal dressing are not already included in the baseline thermal free-energy $V_{T}$ used in the numerics.
@@ 2.3 Phase transition parameters
For analytic intuition, one can approximate the effective potential near the origin as $V_{\mathrm{eff}}\!\approx\!\tfrac{1}{2}m_{\mathrm{eff}}^{2}(T)\phi^{2}-ET\phi^{3}+\tfrac{1}{4}\lambda\phi^{4}$, where $E$ collects cubic thermal contributions. A negative $p$ increases the magnitude of $m_{\mathrm{eff}}^{2}(T)$ at low $T$. Thereby deepening the broken minimum and lowering $T_{c}$ and $T_{n}$. This qualitative behavior explains the trends seen in Fig. 2.
@@
2.4 Real-Singlet Extension as a Concrete UV Realization
To illustrate how a temperature-dependent quadratic correction of the form $c\,T^{p}\phi^{2}$ naturally emerges in a realistic particle-physics model, we briefly summarize the relevant features of the real-singlet extension of the Standard Model (xSM). This minimal scalar extension is well known to support a strong first-order electroweak phase transition through a tree-level barrier, while simultaneously generating thermal self-energy corrections to the Higgs field that match the structure of Eq. (1).
Lagrangian and mixing structure
The scalar potential of the $Z_{2}$-symmetric xSM is
$V(H,S)=-\mu_{H}^{2}|H|^{2}+\lambda_{H}|H|^{4}+\frac{1}{2}\mu_{S}^{2}S^{2}+\frac{1}{4}\lambda_{S}S^{4}+\frac{1}{2}\lambda_{HS}|H|^{2}S^{2},$ (14)
where $H$ is the Higgs doublet and $S$ is a real gauge singlet. After electroweak symmetry breaking, the fields mix due to the portal interaction $\lambda_{HS}|H|^{2}S^{2}$, producing the mass eigenstates $h_{1}$ and $h_{2}$ with mixing angle $\theta$.
Thermal self-energy and mapping to $c\,T^{p},\phi^{2}$
The one-loop finite-temperature Higgs self-energy receives a contribution from singlet loops,
$\Pi_{h}(T)\;\supset\;\frac{\lambda_{HS}}{12}\,T^{2}\;F(m_{S}/T),$ (15)
where $F(x)$ captures the deviation from the pure high-temperature limit and interpolates between $F(x)\!\to\!1$ for $x\ll 1$ and $F(x)\!\propto\!xe^{-x}$ for $x\gg 1$. In the regime where the singlet is non-relativistic during the phase transition, $m_{S}\gtrsim T$, the Boltzmann suppression produces a temperature dependence that is well approximated by a power law,
$F(m_{S}/T)\;\simeq\;A\,T^{\,p},\qquad p<0,$ (16)
with the exponent determined by the dominant microscopic scattering channels. Inserting this into Eq. (15) and using the local approximation for the one-particle-irreducible effective potential yields
$\Delta V_{\text{xSM}}(\phi,T)=\Pi_{h}(T)\,\frac{\phi^{2}}{2}\simeq c\,T^{p}\phi^{2},$ (17)
where the matching coefficient is
$c\;=\;\frac{\lambda_{HS}}{24}\,A.$ (18)
Thus the xSM provides a concrete UV origin for both the sign and the magnitude of the $(c,p)$ parameters used in our phenomenological analysis.
Phase-transition relevance
The portal interaction $\lambda_{HS}$ also generates a tree-level barrier in the scalar potential when the singlet acquires a thermal expectation value or develops temperature-dependent mass corrections. Numerous studies have shown that for natural parameter choices, the xSM supports a strongly first-order electroweak phase transition consistent with Higgs-coupling constraints. In such regions of parameter space, the thermal self-energy term in Eq. (17) constitutes an additional, dynamically evolving contribution to the curvature of the effective potential, and therefore fits squarely into the general framework developed in this paper.
This example demonstrates explicitly that the parametrization in Eq. (1) is not merely phenomenological, but directly realizable in well-motivated extensions of the Standard Model. In the remainder of this work, we focus on the $(c,p)$ description, which efficiently captures the leading impact of such thermal dressing effects on phase-transition dynamics and the resulting gravitational-wave signal.
We emphasize that the assumption $m_{S}\gtrsim T$ does not conflict with the presence of a tree-level barrier or the realization of a first-order phase transition. In singlet-extended scalar theories, the barrier responsible for the phase transition originates from the zero-temperature structure of the scalar potential and, where relevant, from light degrees of freedom whose thermal functions admit a high-temperature expansion. In contrast, the singlet mediator $S$ responsible for the temperature-dependent mass correction $\delta m_{\phi}^{2}(T)$ need not be light and is treated separately through its static thermal self-energy. The heavy-mediator regime $m_{S}\gtrsim T$ therefore applies only to the sector generating the effective $c\,T^{p}\phi^{2}$ contribution and does not affect the validity of the tree-level barrier or the phase transition dynamics. Accordingly, no assumption of a global high-temperature expansion is made, and different thermal regimes are consistently applied to different sectors of the theory.